27% of 350= 12% of 8= 7% of 350=
Answer by stanbon(75887)

27% of 350= (27/100)*350 = 27*3.5= 94.5
------------
12% of 8= (12/100)*8 = (3/25)*8= 24/25= 0.96
-----------
7% of 350= (7/100)*350 = 7 * 3.5 = 24.5
Cheers,
Stan H.